Circular No. 6305

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

(2060) CHIRON H. Campins writes regarding near-infrared observations with M. Rieke,

using the University of Arizona’s 2.3-m reflector (+ 256×256 HgCdTe array camera NICMOS3; range 1–2.5 μm, pixel size ∼ 0′′.6): “We estimate for (2060) Chiron a 2.2-μm magnitude of K = 14.24 ± 0.05 within a 3′′.5diameter aperture at ∼ Feb. 8.35 UT (∼ 6 days before perihelion). It is remarkable that no coma is apparent in our images; a preliminary reduction indicates that the radial profiles of the Chiron images do not differ from those of two nearby stars of similar magnitude. Any coma, if present, is at most a few percent the brightness of the nucleus. We are in the process of obtaining a rotational lightcurve of the nucleus.”

RX J0117.6–7330 AND RX J0550.0–7151 P. A. Charles and K. A. Southwell, Oxford University; and D. O’Dono-

ghue, University of Cape Town, report: “We observed the 14th-magnitude star close to the Small Magellanic Cloud (SMC) x-ray nova RX J0117.6– 7330 (IAUC 6282) with the Anglo-Australian Telescope (AAT) on Jan. 18 UT and monitored it with the South African Astronomical Observatory (SAAO) 1.9-m telescope from Jan. 24 to 28. The AAT spectrum revealed strong, narrow (250 km/s) Balmer emission (Hα, EW = 1.3 nm; Hβ, EW = 0.13 nm) superposed on a blue continuum with weak He I and C III/N III absorption, supporting its identification as the optical counterpart. With a magnitude of V = 14.2, we derive M_{v} = −4.6, which is consistent with a massive OB star in the SMC. No photometric variations greater than 0.05 mag were seen during the SAAO run.

“Our AAT spectra of the red star close to the LMC supersoft source RX J0550.0–7151 (IAUC 6278) reveal strong, narrow (160 km/s) Balmer emission superposed on the spectrum of a cool star, thus confirming its symbiotic nature.”

COMET C/1995 O1 (HALE-BOPP) Total visual magnitude estimate (cf. IAUC 6273) by T. Lovejoy, Jim-

boomba, Queensland, Australia (0.25-m reflector): Feb. 2.75 UT, 8.8 (coma diameter 4′; 8′ tail).

1996 February 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green